How much pure water must bemixed with 6 pints of 60% developer to produce a mixture that is 16% developer?
Answer by robertb(5830)

0.60*6 + 0.00x = 0.16(6 + x)
==> 3.6 = 0.96 + 0.16x
==> 2.64 =0.16x
==> x = 16.5 pints of pure water.